OpenStudy (anonymous):

Hey ! What is the eqn of the tangent line ((x-2)^2)/4+(y+1)^2=1 @ point (3,-(2+sqrt3)/2)? i think the derivative of the eqn is (-1/2x+1)/2y+2.Helpp:))

OpenStudy (anonymous):

is this \[\frac{(x-2)^2}{4+(y+1)^2}=1\]?

OpenStudy (anonymous):

if so first step would be to write \[(x-2)^2=4+(y+1)^2\]

OpenStudy (anonymous):

or maybe even \[(x-2)^2-(y+1)^2=4\]

OpenStudy (anonymous):

noo..almost..the 1st one was more right..excep tht the (y+1)^2 was added

OpenStudy (anonymous):

Satellite:its an ellipse :P That'll show you how to fix the equation.

OpenStudy (anonymous):

\[(x-2)^{2}/4 + (y+1)^{2}\]

OpenStudy (anonymous):

Satellite:its an ellipse :P That'll show you how to fix the equation.

OpenStudy (anonymous):

ok let me try again. \[\frac{(x-2)^2}{4}+(y+1)^2=1\]?

OpenStudy (anonymous):

yess:))

OpenStudy (anonymous):

ooh yes @malevolence got it i am a little slow. other one was hyperbola i think

OpenStudy (anonymous):

Yeah haha

OpenStudy (anonymous):

ok so taking derivatives just as easy we get \[\frac{x-2}{2}+2(y+1)y'=0\]

OpenStudy (anonymous):

lets see if i can solve this for y' without messing up. i get \[y'=\frac{-x+2}{4(y+1)}\] hows that look?

OpenStudy (anonymous):

good!:)

OpenStudy (anonymous):

whew i messed up on the last one ok now we plug an chug yes?

OpenStudy (anonymous):

i guess so..thts wer im stuck:/

OpenStudy (anonymous):

oh well it looks pretty annoying. \[x=3\] \[y=-\frac{2+\sqrt{3}}{2}\]?

OpenStudy (anonymous):

maybe not so bad. numerator of our fraction is 5 yes?

OpenStudy (anonymous):

ok it is -1 fine

OpenStudy (anonymous):

numerator is -1. now lets look at \[y=-\frac{2+\sqrt{3}}{2}=-1-\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

first we add 1 to get \[y+1=-\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

the - gets dsitributed in the 2+ sqrt 3..so the y coordinate is (-2-sqrt3 )/2 right

OpenStudy (anonymous):

then we multiply by 4 to get \[-2\sqrt{3}\]

OpenStudy (anonymous):

yes if it is the way you wrote it we distribute the - sign

OpenStudy (anonymous):

yeah:)

OpenStudy (anonymous):

so after all is said and one, after we add 1, multiply by 4 we get the denominator is just \[-2\sqrt{3}\] and so the whole fraction is \[\frac{1}{2\sqrt{3}}\]

OpenStudy (anonymous):

assuming i did not make another arithmetic mistake

OpenStudy (anonymous):

@malevolence how'd i do?

OpenStudy (anonymous):

i got..\[-1\div -4-2\sqrt{3}\]

OpenStudy (anonymous):

if i plug in the x and y coordiante given

OpenStudy (anonymous):

course you are not done yet. you have to find the equation of the line. use the almighty point-slope formula to get \[y+\frac{2+\sqrt{3}}{2}=\frac{1}{2\sqrt{3}}(x-3)\] etc

OpenStudy (anonymous):

@sanhita that is not what i got

OpenStudy (anonymous):

sry..minez was wrong..

OpenStudy (anonymous):

\[y=-\frac{2+\sqrt{3}}{2}\] yes

OpenStudy (anonymous):

i actually got -1/2-sqrt 3

OpenStudy (anonymous):

lets take it one step at a time

OpenStudy (anonymous):

y=(-2-sqrt 3)/2...when we distribute the -

OpenStudy (anonymous):

which is also \[-1-\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now we add 1 yes?

OpenStudy (anonymous):

denominator is \[4(y+1)\]

OpenStudy (anonymous):

yes..n the numerator wud b -2-sqrt 3?

OpenStudy (anonymous):

first we need to add 1. then multiply by 4

OpenStudy (anonymous):

yea..aftr doin tht...thts wat i got

OpenStudy (anonymous):

sry not -2..2-sqrt 3

OpenStudy (anonymous):

ok one step at a time \[y=-1-\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

\[y+1=-1-\frac{\sqrt{3}}{2}+1=-\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

yup:)

OpenStudy (anonymous):

and finally \[4(y+1)=-2\sqrt{3}\]

OpenStudy (anonymous):

ohhyeaa..rightt

OpenStudy (anonymous):

and finally finally the whole fraction is \[\frac{1}{2\sqrt{3}}\]

OpenStudy (anonymous):

haha yes..sry abt the confusion:)

OpenStudy (anonymous):

and finally finally finally you have to find the equation of the line. i started it somewhere above, i will get you finish it

OpenStudy (anonymous):

well this is wer i am at so far...y+1+sqrt3/2=(1/(2*sqrt3))x-3/(2*sqrt3)

OpenStudy (anonymous):

simplified it is...\[y=1x/\left( 2\sqrt{3} \right)-3/(2\sqrt{3})-1\]

OpenStudy (anonymous):

thnks u soo much 4 ur work n patience!:)

OpenStudy (anonymous):

looks like you got it. i am going to assume you did that last bit of arithmetic correctly. good work!